OpenStudy (anonymous):

Can someone give me the answer plss! Drey makes $10 an hour plus $15 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. PART A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. PART B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. PART C: Drey earned $490 in 1 week. How many hours (regular plus overtime) did he work?

OpenStudy (lynfran):

x=10

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

im just asking because i think its the same thing & someone told me it wasnt 10

OpenStudy (lynfran):

\[-\frac{ 1 }{ 2 }x+2=-x+7\]

OpenStudy (lynfran):

thats the question right

OpenStudy (anonymous):

correct

OpenStudy (lynfran):

so if we substitute x=10 on the right hand side of the equation we get -10+7=-3.....and on the left hand side we get -5+2=-3 ...so the is true x=10

OpenStudy (anonymous):

ok, can you help me with one more? its a little bit lengthy

OpenStudy (lynfran):

sure

OpenStudy (anonymous):

Drey makes $10 an hour plus $15 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. Part B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. Part C: Drey earned $490 in 1 week. How many hours (regular plus overtime) did he work?

OpenStudy (lynfran):

did the question state anything about how many hrs per day drey work?

OpenStudy (anonymous):

no thats all it says

OpenStudy (lynfran):

@DullJackel09

OpenStudy (lynfran):

im thinking the equation for part A would be w=xd and d would equals 7 since 7day in a week, x would be the #of hours and w be the wages

OpenStudy (lynfran):

@ganeshie8